OpenStudy (anonymous):

What is the empirical formula of a compound containing 24.56% potassium, 34.81% manganese, and 40.50% oxygen? KMnO KMnO2 K0.63Mn0.63O2.53 KMnO4 K3MnO2

OpenStudy (anonymous):

Assume a 100g sample to make things easier; so 24.56% of 100g is 24.56g, for example. The empirical formula is simply the ratio of the moles of each substance. So convert each to moles so that you can compare them: 24.56g K *(1 mol/39.10g)= .6281 mol K 34.81g Mn *(1mol/54.94)= .6336 mol Mn 40.50g O *(1mol/16.00g)= 2.531 mol O Remember that we are trying to find the ratio between them. The easiest way to do that is to divide each by the smallest number: .6281 mol K /.6281 mol = 1 .6336 mol Mn /.6281mol = 1.0 2.531 mol O /.6281 mol =~ 4.0 So you have a ratio of 1:1:4 (there is some degree of approximation that you have to make) So the empirical formula is: KMnO4 (the 4 is a subscript)

OpenStudy (anonymous):

%'s are % by weight, so use them as grams, & use their molar masses, i norder to find moles: (24.7 grams of K) / (39.1 grams per mol K) = 0.63 mol K (34.8grams of Mn) / (54.9 grams per mol Mn) = 0.63 mol Mn (40.5 grams of O) / (16 grams per mol O) = 2.53 mol O divide by the smallest number of moles, in order to ratio the moles: (0.63 mol K) / (0.63) = 1 mol K (0.63 mol Mn) / (0.63) = 1 mol Mn (2.53 mol O) / (0.63) = 4 mol O your answer is KMnO4